Grade 12Mechanics

A satellite of mass m initially at rest on earth is launched into a circular orbit at a height equal to the radius of the earth the minimum energy required is

Samyra Singh
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
G = Gravitational Constant. M : mass of Earth m : mass of object. r : radius of earth , h = height of orbit
it is based on energy conservation.

 
Ui = -GMm/r

 
Uf = -GMm/(r+h)  ; h=r

 
so, energy required =difference in potential energy

 
                                = -GMm/2r + GMm/r =GMm/2r